# exampleframelinecolor=gray!30, linewidth=0.5pt, backgroundcolor=gray!5, roundcorner=2pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt, skipabove=skipbelow=

### v1

@@ 4 Can Harmful Subspaces Be Removed?
Having analyzed helpful and harmful updates in isolation, we now consider a more realistic scenario: contaminated FT. This involves adding a small fraction of harmful examples to an otherwise benign dataset, producing updates that blend both signals. Contaminated data is particularly dangerous because it can degrade alignment without obvious signs. Prior work shows that even limited contamination can erode safety, causing models to revert to unsafe behaviors. While earlier experiments identified expressive subspaces, we now ask the reverse: can we remove harmful components from an update? We test whether filtering specific subspaces, particularly those aligned with the dominant directions of the alignment matrix, can reduce harmfulness while preserving utility.
@@
We construct a contaminated dataset by mixing 20% harmful data from BeaverTails with 80% of the 20K MetaMathQA subset. FT on this mixture yields a single contaminated update, $\Delta_{T}$. To suppress harmful behavior, we apply the orthogonal projection strategy from Section 2, removing components along the top-$k$ alignment directions. Specifically, we compute $\tilde{\Delta}_{T}=P_{k}^{\perp}\Delta_{T}$, where $P_{k}^{\perp}$ projects onto the complement of the alignment subspace. We evaluate the resulting models on GSM8K (utility) and AdvBench (harmfulness). Our goal is to test whether removing alignment-aligned components can reduce harmfulness while preserving task performance.
@@
Figure 3 shows the effects of orthogonal projection on retained energy, utility, and harmfulness. As $k$ increases, implying more of the update is removed, the retained energy declines steadily across all projection types (random, top-$k$, and random-$k$). Utility and harmfulness scores (Figure 3, Table 2) follow a similar downward trend. However, the rate of decline differs by projection strategy. Removing top-$k$ alignment components reduces utility more sharply than random projections. At the same time, harmfulness decreases at a similar rate, indicating no selective suppression of harmful behavior. In effect, safety improvements come at a proportional cost to task performance, with no clear advantage in targeting the alignment subspace. We present results on all models in Table 4 (Appendix C).
@@
These results indicate that the alignment subspace does not uniquely encode safety or harmfulness but rather captures directions broadly important for learning. Removing these directions degrades both utility and harmfulness at similar rates. If harmful behavior were confined to distinct subspaces, we would expect a steeper drop in harmfulness than utility, yet this is not observed. Even if safety-relevant directions exist, they are not recoverable from the alignment matrix alone, especially under contamination. The update blends helpful and harmful objectives, making its projection agnostic to intent. As a result, orthogonal projection fails to selectively suppress harmful behavior. Subspace filtering based on alignment directions imposes a strict tradeoff: gains in safety come with proportional utility loss. This challenges the effectiveness of subspace-based defenses under contaminated FT.
@@ 5 Are Safety Weight Subspaces Distinct?
Figure 4: Mode Subspace Overlap (MSO) at the 70- and 85- percentile layers for pairwise comparisons of the dominant subspaces from Harmful fine-tuned (H), Instruction-tuned (I), and Base (B) models.
A natural question is whether a dedicated region of parameter space, what we might call a safety subspace, captures safety-specific behavior. Such a subspace should meet two criteria: (i) safety-relevant updates, whether from alignment or harmful FT, should lie significantly within it; and (ii) task-specific updates unrelated to safety should have minimal overlap, with projections onto the subspace leaving model safety unchanged. Crucially, these properties must generalize across tasks and datasets to rule out dataset-specific artifacts. Our earlier results argue against the top subspaces of the alignment matrix meeting these criteria. These directions are highly sensitive to any update, helpful or harmful, but do not isolate safety. Still, it remains open whether some other set of directions, possibly outside the alignment subspace, could fulfill this role. To explore this, we directly compare the dominant subspaces of different update types.
@@
We compare the principal subspaces of 3 updates: the alignment update $\Delta_{A}$ (from base to aligned model), the harmful FT update $\Delta_{\mathrm{T}}^{\text{Harmful}}$ (trained on BeaverTails), and the useful update $\Delta_{\mathrm{T}}^{\text{Useful}}$ (trained on a 20K subset of MetaMathQA). Notably, the negated alignment update, $-\Delta_{A}$, reverses alignment by pushing the model back toward its unaligned base state, effectively acting as a harmful update and serving as a useful reference. For a given energy threshold $\eta\in(0,1]$, we compute $\mathrm{MSO}(\cdot,\cdot;\eta)$ (Section 2 for three pairs: (i) $\big{(}\Delta_{\mathrm{T}}^{\text{Useful}},\Delta_{\mathrm{T}}^{\text{Harmful}}\big{)}$, to assess whether helpful and harmful FT affect similar subspaces; (ii) $\big{(}\Delta_{\mathrm{T}}^{\text{Useful}},-\Delta_{A}\big{)}$, to test alignment between helpful updates and reversed alignment; and (iii) $\big{(}\Delta_{\mathrm{T}}^{\text{Harmful}},-\Delta_{A}\big{)}$, to compare two harmful directions. We sweep over $\eta$, with small values isolating high-energy directions and larger values approaching full-rank overlap. We include the random-subspace baseline $\max(k_{V},k_{W})/d$; values above this baseline indicate significant geometric alignment, while values near it suggest chance-level overlap.
@@
Figure 4 shows the pairwise overlap between the dominant subspaces (top-$k$ directions) of each update. All pairs exhibit greater overlap than random baselines, indicating shared structure. However, the strongest overlap is between the useful and harmful updates, not between alignment and harmful updates, as one might expect if safety were a shared component. This is a key finding. If a safety subspace existed, it would likely appear in the shared directions between alignment and harmful updates, which affect safety in opposite ways. This lack of substantial overlap suggests that no consistent, linear safety-specific subspace exists.
@@
Together with earlier results, these findings suggest that safety-relevant updates do not lie in a well-defined or isolatable subspace. Instead, both alignment and harmfulness operate over complex, task-dependent, and likely non-linear directions. The high overlap between harmful and helpful update subspaces supports our earlier hypothesis: these directions form a general learning subspace, expressive across tasks but agnostic to safety. We find no evidence for a distinct safety subspace. Updates that influence safety, positively or negatively, do not share dominant directions. Any shared structure reflects general learning capacity rather than safety-specific behavior. As such, geometric separation of alignment remains elusive, and linear subspace methods cannot cleanly isolate safety in parameter space.
6 Do Safety Subspaces Exist In Representation Space?
@@
So far, our analysis has focused on the weight space, probing whether certain update directions correspond to safety-related behavior. Across experiments, we found no evidence for distinct subspaces encoding safety at the parameter level. However, safety may instead manifest through how inputs interact with the model, specifically, through the regions of representation space they activate. This motivates a final question: do safety-relevant inputs elicit distinct activation patterns, even if their corresponding weight updates overlap? While weight updates may distribute energy broadly, inputs could selectively activate specific directions. This perspective also offers a possible explanation for earlier results: even low-energy projections onto alignment directions produced strong behavioral effects, likely because inputs activated those directions disproportionately.
@@
We compare internal activations induced by different prompt categories. Specifically, we pass useful (benign) prompts from the MATH dataset [17] and harmful prompts from BeaverTails (test set) and ToxiGen [14] through three models: the aligned model, the useful fine-tuned model, and the contaminated fine-tuned model. For each prompt, we record the hidden state of the last generated token at each transformer layer $\ell\in{0,\dots,L}$. At each layer, these hidden states are stacked into activation matrices of shape $\mathbb{R}^{n\times d}$, where $d$ is the model’s hidden size and $n$ is the number of prompts (5000 for each dataset). We compute MSO (see Section 2) between activation matrices corresponding to the prompts from different datasets, sweeping over energy thresholds $\eta$. Lower values of $\eta$ capture high-energy activation modes, while higher values approximate full-rank comparisons. We plot MSO curves alongside the random-subspace baseline $\max(k_{\text{Useful}},k_{\text{Harmful}})/d$, and report averages over layers in the 65–90 % depth percentile.
@@
Figure 5 reports MSO values across all pairs of prompt categories. Useful and harmful prompts consistently exhibit overlap above the random baseline, indicating activation of shared high-energy subspaces in representation space. Notably, the overlap between the two harmful prompt sets is not consistently higher than their overlap with helpful prompts; in some cases, the useful–harmful overlap is greater than the harmful–harmful one. The degree of overlap also varies across model configurations. Some models show strong alignment even in the top subspaces, while others exhibit more gradual increases in overlap, becoming significant only at higher energy thresholds. This variability suggests that representational similarity is influenced more by model-specific factors than by the safety content of the prompts alone. Results on more models are provided in Figure 6 (Appendix D).
@@
These observations suggest that while all prompt types activate shared subspaces more than expected by chance, there is no evidence of a distinct safety-violating subspace. If such a subspace existed, activations from harmful prompts would consistently exhibit greater mutual overlap than with useful prompts, which is not the case. Instead, the results indicate that prompts with differing safety implications are processed through broadly overlapping representations. This supports our earlier hypothesis: the directions most responsible for driving behavior reflect general-purpose representational subspaces rather than safety-specific ones. These directions are activated across tasks and prompt types, implying that LLMs do not internally separate “safe” and “unsafe” activation modes, but instead rely on shared, high-impact subspaces. We find no evidence of a distinct safety subspace in representation space. Useful and harmful prompts show substantial overlap, even across prompt sets with very different behavioral consequences. Combined with our findings in weight space, these results suggest that both aligned and harmful behaviors emerge from shared representational mechanisms rather than separable subspaces.
@@ 7 Conclusion
This work set out to investigate how safety alignment is encoded in LLMs, and whether it can be isolated through geometric structure in weight or activation space. Motivated by the challenge of preserving alignment under continued fine-tuning, particularly in adversarial or contaminated settings, we conducted a systematic study across four experiments and five open-source LLMs, examining both parameter updates and internal representations. Our findings challenge the common assumption that alignment corresponds to safety-specific subspaces. Subspaces with high behavioral impact are not unique to alignment; they enhance both utility and harmfulness, and their removal degrades both. This indicates that these directions reflect general-purpose learning rather than safety alone. Moreover, harmful and helpful prompts activate overlapping regions of representation space, offering no evidence for distinct “safety activation” geometry. Together, these results suggest that safety alignment is not cleanly separable in geometric terms. While this complicates subspace-based defenses, it also highlights the potential of high-impact directions, if appropriately constrained, for guiding both safe fine-tuning and activation-level control. More broadly, our work calls for rethinking geometric assumptions in interpretability and alignment, and for developing methods that engage with the entangled nature of learned representations.
8 Limitations
Our study focuses on linear subspaces, providing a principled first step toward understanding the geometric structure of safety alignment. While we do not explore non-linear representations, our framework could be extended in future work to capture richer geometric phenomena. Due to computational constraints, we conduct experiments on small and mid-sized LLMs with publicly available base and instruction-tuned variants. These models offer a controlled and interpretable setting, though generalization to larger architectures and alternative alignment methods remains an important direction for future investigation.
@@ Appendix A Related Work
Large Language Models (LLMs) do not inherently follow instructions and often exhibit socially undesirable behaviors. To address this, various post-training methods, instruction-tuning and reinforcement learning from human feedback, are applied to align base LLMs with human values and improve their instruction-following capabilities [37, 46, 55, 44]. However, studies have shown that fine-tuning these aligned models on harmful data can undo this alignment, restoring their original, socially unacceptable behaviors [59]. This unalignment phenomenon has been demonstrated in both open-source models [60, 29] and proprietary models [42, 4, 64] via publicly available fine-tuning APIs, thereby exposing a new attack surface [23, 9, 27]. Moreover, even fine-tuning on benign downstream tasks can degrade alignment [16, 15].
@@
To safeguard aligned LLMs against unalignment during fine-tuning, defenses have been proposed at three stages of the pipeline: the alignment stage, the fine-tuning stage, and the post-processing stage. The effectiveness of these defense methods is evaluated using downstream model utility and harmfulness [20].
@@
Alignment stage defenses update the initial instruction-tuning process to ensure that downstream fine-tuning cannot easily overwrite the model’s safety behavior. One approach augments the alignment loss, making harmful representations harder to recover during fine-tuning updates [45]. Another line of work relies on safety-oriented data curation to preserve alignment under downstream fine-tuning[33]. Adversarial and meta-learning techniques have also been combined to develop tamper-resistant methods that prevent harmfulness while maintaining task performance [48]. A separate strategy introduces a regularization term to the alignment loss, which has been shown to preserve safety after fine-tuning [22]. Perturbing safety-critical layers during instruction-tuning has also been shown to protect alignment [32]. Additional work traces unalignment to excessive dependence on maximum-likelihood training, motivating an integrity preserving variant of this method [6]. A study on “shallow alignment” also shows that instruction-tuning influences only the first few output tokens, whereas deeper alignment improves robustness [41].
@@
Fine-tuning stage defenses modify the fine-tuning process to ensure that the model’s alignment is preserved after update. One class of defenses focuses on data curation, augmenting the fine-tuning dataset to maintain alignment after update [5, 13]. Another approach uses safety examples prefixed with a secret prompt, which act as backdoor triggers to reactivate safe behavior after fine-tuning [52]. A data ranking based strategy has also been proposed, where low-quality data is down-ranked and high-quality data is up-ranked to better preserve safety [47]. It has also been shown that prompt templates play an important role; removing the safety prompt during fine-tuning and reintroducing it at inference time can maintain alignment [35].
Optimization based defenses are another type of fine-tuning stage defenses. One line of work splits fine-tuning into an alignment phase and a utility phase, safeguarding both safety and task performance [21]. Another approach combines safety and helpfulness objectives into a single loss [65].
Parameter level methods can also be used to preserve safety. One strategy identifies safety neurons and updates only those parameters [66]. Another approach involves localizing safety layers and freezing their gradients, which has been shown to prevent unalignment [31]. Another line of work explores constraining parameter changes to directions orthogonal to existing safety features, showing that this method preserves alignment [30]. It has also been shown that harmful data can be filtered by matching fine-tuning embeddings against the top-k singular vectors of an activation matrix generated using a harmful dataset [7].
@@
Post-processing stage defenses adjust the fine-tuned model to restore alignment and preserve usefulness. One approach adds a safety vector, defined as the difference between aligned and unaligned weights, to the fine-tuned parameters to regain safe behavior [3]. Another line of work projects the fine-tuning update onto the alignment vector when their similarity drops below a threshold, or selectively merges layers from the fine-tuned and aligned models under the same criterion to achieve a similar effect [18, 10]. A third strategy removes parameters identified as harmful after fine-tuning to restore alignment [19]. It has been shown that safety directions in attention-head activations can also be located and used for targeted intervention [67] to realign the fine-tuned model. Another method detects update parameters whose signs contradict the original alignment and removes them [57]. Additional work restores safety-critical neurons [61], fuses aligned and fine-tuned models [62], or adds an optimized post-hoc perturbation to recover alignment [53].
@@
Recent studies have examined how LLMs express safety over neurons, layers, and activations. One study finds that safety related information is language agnostic, identifies parameters whose modification affects alignment, and shows that freezing these parameters during fine-tuning does not ensure safety [40]. Another line of work locates sparse regions in parameter space whose removal weakens alignment, and likewise observes that freezing these regions alone is insufficient to maintain model alignment [54]. A separate analysis maps a safety basin in weight space, noting that random perturbations inside the basin leave safety intact, whereas fine-tuning moves weights outside it [39]. Finally, work on the activation residual stream isolates a refusal direction, removing this direction prevents refusal to harmful prompts, while adding it triggers refusal to benign ones [2].
Appendix B Do Alignment Subspaces Encode Safety?
We provide additional results in Table 3 to support the analysis presented in Section 3.

### v2

@@ 4 Can Harmful Subspaces Be Removed?
Having analyzed helpful and harmful updates in isolation, we now turn to a more realistic scenario: contaminated FT. This setting involves adding a small fraction of harmful examples to an otherwise benign dataset, producing updates that blend both signals. Prior work has shown that even limited contamination can erode safety, causing models to revert to unsafe behaviors (64, 32, 46, 4, 68). While earlier experiments identified expressive subspaces, we now ask the reverse question: can harmful components of an update be removed? We test whether filtering specific subspaces, particularly those aligned with the dominant directions of the alignment or safety-tuned matrix, can reduce harmfulness while preserving utility.
@@
We construct a contaminated dataset by mixing 20% harmful data from BeaverTails with 80% of the 20K MetaMathQA subset. FT on this mixture produces a single contaminated update, $\Delta_{T}$. To suppress harmful behavior, we apply the orthogonal projection strategy from Section 2, removing components along the top-$k$ alignment directions. Specifically, we compute $\tilde{\Delta}_{T}=P_{k}^{\perp}\Delta_{T}$, where $P_{k}^{\perp}$ projects onto the complement of the alignment subspace. We then evaluate the resulting models on GSM8K (utility) and AdvBench (harmfulness). Our objective is to test whether removing alignment-aligned components can reduce harmfulness while preserving task performance.
@@
Figures 3 and 7 (Appendix D) show the effects of orthogonal projection on retained energy, utility, and harmfulness. As $k$ increases, meaning more of the update is removed, retained energy declines steadily across all projection types (random, top-$k$, and random-$k$). Utility and harmfulness scores (Figure 3, Table 2) follow a similar downward trend. The rate of decline, however, differs by projection strategy. Removing top-$k$ alignment components reduces utility more sharply than random projections, while harmfulness decreases at a similar rate. This indicates no selective suppression of harmful behavior. In effect, safety gains come at a proportional cost to task performance, with no clear advantage in targeting the alignment subspace. Comprehensive results for all models are provided in Table 5 (Appendix D).
@@
These results establish that the top subspaces of alignment or safety-tuned updates do not uniquely encode safety or harmfulness. Removing these directions degrades both utility and harmfulness at similar rates. If harmful behavior were confined to distinct subspaces, we would expect a steeper drop in harmfulness than in utility, yet this is not observed. Even if safety-relevant directions exist, they cannot be recovered from the alignment or safety-tuned matrices alone, particularly under contamination. The update blends helpful and harmful objectives, making its projection agnostic to intent. As a result, orthogonal projection fails to selectively suppress harmful behavior. Thus, subspace filtering based on alignment directions imposes a strict trade-off: improvements in safety come only at a proportional cost to utility.
@@ 5 Are Safety Weight Subspaces Distinct?
Figure 4: Mode Subspace Overlap (MSO) at the 70- and 85- percentile layers for pairwise comparisons of the dominant subspaces from Harmful fine-tuned (H), Aligned (A), and Base (B) models.
A natural question is whether a dedicated region of parameter space, which we might call a “safety subspace,” captures safety-specific behavior. Such a subspace should satisfy two criteria: (i) safety-relevant updates, whether from alignment or harmful FT, should lie predominantly within it; and (ii) task-specific updates unrelated to safety should have minimal overlap, with projections onto the subspace leaving model safety unchanged. Our earlier results argue against the top subspaces of the alignment or safety-tuned matrices meeting these criteria. Nevertheless, it remains open whether some other set of directions, possibly outside these subspaces, could fulfill this role. To investigate this, we directly compare the dominant subspaces of different update types.
@@
We compare the principal subspaces of three updates: the alignment update $\Delta_{A}$ (from the base to the aligned model), the harmful FT update $\Delta_{\mathrm{T}}^{\text{Harmful}}$ (trained on BeaverTails), and the useful FT update $\Delta_{\mathrm{T}}^{\text{Useful}}$ (trained on a 20K subset of MetaMathQA). Notably, the negated alignment update $\Delta_{A}$ reverses alignment by pushing the model back toward its unaligned base state, effectively acting as a harmful update and serving as a useful reference point. We repeat these experiments for safety-tuned updates as well. For a given energy threshold $\eta\in(0,1]$, we compute $\mathrm{MSO}(\cdot,\cdot;\eta)$ (Section 2) for three pairs: (i) $\big(\Delta_{\mathrm{T}}^{\text{Useful}},\Delta_{\mathrm{T}}^{\text{Harmful}}\big)$, to assess whether helpful and harmful FT affect similar subspaces; (ii) $\big(\Delta_{\mathrm{T}}^{\text{Useful}},-\Delta_{A}\big)$, to test the relationship between helpful updates and reversed alignment; and (iii) $\big(\Delta_{\mathrm{T}}^{\text{Harmful}},-\Delta_{A}\big)$, to compare two harmful directions. We sweep over $\eta$, with smaller values isolating high-energy directions and larger values approaching full-rank overlap. As a baseline, we include the random-subspace expectation $\max(k_{V},k_{W})/d$; values above this baseline indicate significant geometric alignment, while values near it suggest chance-level overlap.
@@
Figures 4 and 8 (Appendix E) show the pairwise overlap between the dominant subspaces (top-$k$ directions) of each update. All pairs exhibit greater overlap than random baselines, indicating shared structure. However, in Figure 4, the strongest overlap is between the useful and harmful updates, rather than between alignment and harmful updates, as one might expect if safety were a shared component. This is a key finding. If a safety subspace existed, it would likely appear in the shared directions between alignment and harmful updates (or between safety and harmful updates), which affect safety in opposite ways. The absence of such overlap suggests that no consistent, linear safety-specific subspace exists. For safety-tuned models (Figure 8), the strongest overlap occurs between the useful and safety-specific updates, an even more counterintuitive result. This implies that, in terms of subspace overlap, the useful update lies closer to the safety-specific update than the harmful update does. Strikingly, this overlap is much larger than that between the harmful and safety-specific updates, even though, semantically, one might expect the latter to be most similar.
@@
Taken together, our results suggest that safety-relevant updates do not reside in a well-defined or isolatable subspace. Instead, alignment, safety, and harmfulness operate over complex, task-dependent directions. The strong overlap between harmful and helpful update subspaces indicates that these directions form a general learning subspace, expressive across tasks but agnostic to safety. Thus, we find no evidence of a distinct safety subspace, and linear subspace methods cannot cleanly isolate safety in parameter space. This highlights a fundamental limitation in subspace-based defences: attempts to filter safety-relevant components suppress general learning as well.
6 Do Safety Subspaces Exist In Activation Space?
@@
So far, our analysis has focused on weight space, probing whether certain update directions encode safety-related behavior. Finding no evidence of distinct safety subspaces at the parameter level raises a final question: do safety-relevant inputs elicit distinct activation patterns, even when their weight updates overlap? Although weight updates may distribute broadly, inputs might still selectively activate specific directions. We investigate this possibility in the following section.
@@
We compare internal activations induced by different prompt categories. Specifically, we pass useful (benign) prompts from the MATH dataset (18) and harmful prompts from BeaverTails (test set) and ToxiGen (15) through three models: the aligned model, the useful fine-tuned model, and the contaminated fine-tuned model. For each prompt, we record the hidden state of the last generated token (31) at each transformer layer $\ell\in{0,\dots,L}$. At each layer, these hidden states are stacked into activation matrices of shape $\mathbb{R}^{n\times d}$, where $d$ is the hidden size and $n$ is the number of prompts (5000 per dataset). We then compute MSO (see Section 2) between activation matrices from different datasets, sweeping over energy thresholds $\eta$. Smaller values of $\eta$ capture high-energy activation modes, while larger values approximate full-rank comparisons. We plot MSO curves alongside the random-subspace baseline $\max(k_{\text{Useful}},k_{\text{Harmful}})/d$ and report averages over layers in the 65–90% depth percentile.
@@
Figure 5 shows MSO values across all pairs of prompt categories. Useful and harmful prompts consistently exhibit overlap above the random baseline, indicating activation of shared high-energy subspaces in activation space. Interestingly, the overlap between the two harmful prompt sets is not always greater than their overlap with useful prompts; in some cases, the useful–harmful overlap exceeds the harmful–harmful overlap. The degree of overlap also varies across model configurations. Some models show strong alignment even in the top subspaces, while others exhibit more gradual increases, with overlap becoming significant only at higher energy thresholds. This variability suggests that representational similarity depends more on model-specific factors than on the safety content of the prompts. Additional results on other models are provided in Figure 9 (Appendix F).
@@
These observations suggest that while all prompt types activate shared subspaces more than expected by chance, there is no evidence of a distinct safety-violating subspace. If such a subspace existed, activations from harmful prompts would consistently show greater mutual overlap than with useful prompts, which is not observed. Instead, prompts with different safety implications are processed through broadly overlapping representations. This supports our earlier hypothesis: the directions most responsible for behavior correspond to general-purpose representational subspaces rather than safety-specific ones. These directions are activated across tasks and prompt types, indicating that LLMs do not internally separate “safe” and “unsafe” activation modes but instead rely on shared, high-impact subspaces. We therefore find no evidence of a distinct safety subspace even in activation space. Together with our weight-space results, this suggests that both aligned and harmful behaviors arise from shared representational mechanisms rather than separable subspaces.
@@ 7 Conclusion
This work set out to investigate how safety alignment is encoded in LLMs and whether it can be isolated in weight or activation space. Our findings challenge the common assumption that alignment or safety-specific updates correspond to unique “safety subspaces”. Subspaces with strong behavioral impact are not unique to safety; rather, they amplify both utility and harmfulness, indicating that safety is deeply entangled with general learning. Similarly, harmful and useful prompts activate overlapping regions of activation space, providing no evidence for a distinct safety subspace. Together, these results establish that safety alignment is not linearly separable in LLMs. While this complicates the development of subspace-based defenses, it also highlights the potential of high-impact directions, if appropriately constrained, for guiding both safe fine-tuning and activation-level control. More broadly, our work calls for rethinking assumptions in interpretability and alignment research, and for developing methods that explicitly account for the entangled nature of representations.
@@ Appendix A Related Work
Large Language Models (LLMs) do not inherently follow instructions and often exhibit socially undesirable behaviors. To address this, various post-training methods, instruction-tuning and reinforcement learning from human feedback, are applied to align base LLMs with human values and improve their instruction-following capabilities (40, 50, 59, 48). However, studies have shown that fine-tuning these aligned models on harmful data can undo this alignment, restoring their original, socially unacceptable behaviors (63). This unalignment phenomenon has been demonstrated in both open-source models (64, 32) and proprietary models (46, 4, 68) via publicly available fine-tuning APIs, thereby exposing a new attack surface (24, 10, 29). Moreover, even fine-tuning on benign downstream tasks can degrade alignment (17, 16).
@@
To safeguard aligned LLMs against unalignment during fine-tuning, defenses have been proposed at three stages of the pipeline: the alignment stage, the fine-tuning stage, and the post-processing stage. The effectiveness of these defense methods is evaluated using downstream model utility and harmfulness (21).
@@
Alignment stage defenses update the initial instruction-tuning process to ensure that downstream fine-tuning cannot easily overwrite the model’s safety behavior. One approach augments the alignment loss, making harmful representations harder to recover during fine-tuning updates (49). Another line of work relies on safety-oriented data curation to preserve alignment under downstream fine-tuning(36). Adversarial and meta-learning techniques have also been combined to develop tamper-resistant methods that prevent harmfulness while maintaining task performance (52). A separate strategy introduces a regularization term to the alignment loss, which has been shown to preserve safety after fine-tuning (23). Perturbing safety-critical layers during instruction-tuning has also been shown to protect alignment (35). Additional work traces unalignment to excessive dependence on maximum-likelihood training, motivating an integrity preserving variant of this method (7). A study on “shallow alignment” also shows that instruction-tuning influences only the first few output tokens, whereas deeper alignment improves robustness (45).
@@
Fine-tuning stage defenses modify the fine-tuning process to ensure that the model’s alignment is preserved after update. One class of defenses focuses on data curation, augmenting the fine-tuning dataset to maintain alignment after update (5, 14). Another approach uses safety examples prefixed with a secret prompt, which act as backdoor triggers to reactivate safe behavior after fine-tuning (56). A data ranking based strategy has also been proposed, where low-quality data is down-ranked and high-quality data is up-ranked to better preserve safety (51). It has also been shown that prompt templates play an important role; removing the safety prompt during fine-tuning and reintroducing it at inference time can maintain alignment (38).
Optimization based defenses are another type of fine-tuning stage defenses. One line of work splits fine-tuning into an alignment phase and a utility phase, safeguarding both safety and task performance (22). Another approach combines safety and helpfulness objectives into a single loss (69).
Parameter level methods can also be used to preserve safety. One strategy identifies safety neurons and updates only those parameters (70). Another approach involves localizing safety layers and freezing their gradients, which has been shown to prevent unalignment (34). Another line of work explores constraining parameter changes to directions orthogonal to existing safety features, showing that this method preserves alignment (33). It has also been shown that harmful data can be filtered by matching fine-tuning embeddings against the top-k singular vectors of an activation matrix generated using a harmful dataset (8).
@@
Post-processing stage defenses adjust the fine-tuned model to restore alignment and preserve usefulness. One approach adds a safety vector, defined as the difference between aligned and unaligned weights, to the fine-tuned parameters to regain safe behavior (3). Another line of work projects the fine-tuning update onto the alignment vector when their similarity drops below a threshold, or selectively merges layers from the fine-tuned and aligned models under the same criterion to achieve a similar effect (19, 11). A third strategy removes parameters identified as harmful after fine-tuning to restore alignment (20). It has been shown that safety directions in attention-head activations can also be located and used for targeted intervention (71) to realign the fine-tuned model. Another method detects update parameters whose signs contradict the original alignment and removes them (61). Additional work restores/finds safety-critical neurons (65, 6), fuses aligned and fine-tuned models (66), or adds an optimized post-hoc perturbation to recover alignment (57).
@@
Recent studies have examined how LLMs express safety over neurons, layers, and activations. One study finds that safety related information is language agnostic, identifies parameters whose modification affects alignment, and shows that freezing these parameters during fine-tuning does not ensure safety (44). Another line of work locates sparse regions in parameter space whose removal weakens alignment, and likewise observes that freezing these regions alone is insufficient to maintain model alignment (58). A separate analysis maps a safety basin in weight space, noting that random perturbations inside the basin leave safety intact, whereas fine-tuning moves weights outside it (43). Work on the activation residual stream isolates a refusal direction, removing this direction prevents refusal to harmful prompts, while adding it triggers refusal to benign ones (2). Other work shows that safety in models is governed by multiple directions in residual stream space, and that refusal rates drop when prompts avoid tokens activating these directions (41). Finally, a study shows that safety fine-tuning minimally adjusts MLP weights by pushing unsafe inputs into the weights’ null space, leading models to process adversarial prompts as safe (27).
Appendix B Experimental Details
We implemented all experiments using PyTorch (42) and the HuggingFace Transformers library (60). We ran all experiments on a single NVIDIA A6000 GPU (48 GB). To save memory, all base models are initalized in torch.bfloat16 precision. All models are trained using the AdamW optimizer (37). Detailed hyperparameter configurations for full fine-tuning (and safety-tuning) of each model are presented in Table 3.
Optimizer AdamW
Batch size 1
Max. Seq. Len 512
Grad Acc. Steps 32
Epochs 1
Learning Rate $1\times 10^{-5}$
LR Scheduler Cosine
Warmup Ratio 0.02
Appendix C Do Alignment Subspaces Encode Safety?
We provide additional results in Figure 6 and Table 4 to support the analysis presented in Section 3.